andrew is making a birthday card for her sister. he wants to draw flowers as many as the age of her sister. how many flowers he need to draw if her sister is 6 years older than him, and the sum of their ages five years from now is 34
Answer by robertb(5830)

Andrew's age = a
Andrew's sister's age = a+6
Now (a + 5) + (a + 6 + 5) = 34
==> 2a + 16 = 34
==> 2a = 18
==> a = 9, Andrew's age now, and
a + 6 = 15, Andrew's sister's age now.